You have been asked to assess the expected financial impact of each of the following proposals to improve the profitability of credit sales made by your company. Each proposal is independent of the other. Answer all questions. Showing your work may earn you partial credit. Proposal #1 would extend trade credit to some customers that previously have been denied credit because they were considered poor risks. Sales are projected to increase by $200,000 per year if credit is extended to these new customers. Of the new accounts receivable generated, 6% are projected to be uncollectible. Additional collection costs are projected to be 5% of incremental sales, and production and selling costs are projected to be 78% of sales. Your firm expects to pay a total of 30% of its income after expenses in taxes. 1) Compute the incremental income after taxes that would result from these projections:

2) Compute the incremental Return on Sales if these new credit customers are accepted: If the receivable turnover ratio is expected to be 5 to 1 and no other asset buildup is needed to serve the new customers…

3) Compute the additional investment in Accounts Receivable

4) Compute the incremental Return on New Investment

5) If your company requires a 20% Rate of Return on Investment for all proposals, do the numbers suggest that trade credit should be extended to these new customers? Explain.

Proposal #2 would establish local collection centers throughout the region to decrease the time it takes to convert credit payments that are mailed in by check to cash. It is estimated that establishing these collection centers would reduce the average collection time by 2 days (from 5 days to 3 days).

  1. If the company currently averages $20,000 in collections per day, how many dollars will this suggested cash management system frees up?

  1. If all freed up dollars would be used to pay down debt that has an interest rate of 8%, how much money could be saved each year in interest expense?
  1. Do the numbers suggest that this new system should be implemented if its total annual cost is $5200? Explain.
Answer #1
Proposal #1:
1 Incremental income after taxes:
$ $
Incremental sales 200000
Less:
Uncollectible accounts (200000*6%) 12000
Additional collection cost (200000*5%) 10000
Production and selling costs (200000*78%) 156000 178000
Incremental income before taxes 22000
Less: Taxes at 30% 6600
Incremental income after taxes 15400
2 Incremental return on sales=Incremental income after taxes/Incremental sales=15400/200000=0.077=7.7%
3 Additional investment in accounts receivable=Incremental sales/Receivable turnover ratio=200000/5=$ 40000
4 Incremental return on new investment=Incremental income after taxes/Additional investment=15400/40000=0.385=38.5%
5 Yes,Since the Incremental return on new investment is greater than 20%
Proposal #2:
Savings by cash management system=Average collection per day*Reduction in average collection time=20000*2=$ 40000
Savings in interest expense=Savings*8%=40000*8%=$ 3200
No since the cost of implementation ($ 5200) is more than the benefit ($ 3200)